iv) Write Mayer’s formula. The molar heat capacities of a gas at constant pressure
and constant volume are 28.8 J mol−1
K−1
and 20.5 J mol−1
K−1
respectively.
Calculate the gas constant
1
Expert's answer
2018-03-31T09:52:08-0400
The Mayer’s law states
C_(P,m)-C_(V,m)=R
Therefore, the gas constant
R=28.8-20.5=8.3 J/(mol∙K)